f(x)=(x+1)^3 (x-1)^3 prove if its increasing or decreasing using geometrical representation
 

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

(x-1)3(x+1)3= (x2-1)3
on differentiation
f’= 3(x2-1)2.(2x) = +ve for +ve value of x and negative for negative value of x

Hence,it is neither increasing nor decreasing

Thanks